Use two equations in two variables to solve the application.
A rancher raises three times as many cows as horses. If he has 112 animals, how many cows does he have?

Answer by ikleyn(52802)

.
Let C be the number of cows and H be the number of horses.


From the condition, you have this system of 2 equations in 2 unknowns


C = 3H,           (1)
C + H = 112.      (2)


Solve it by the substitution method.
From equation (1), you just have an expression for C.  Substitute it into equation (2), replacing C:


3H + H = 112,

4H = 112

H = 112/4 = 28.


Answer.  28 horses and 3*28 = 84 cows.